# A quasilinear Keller-Segel model with saturated discontinuous advection

### v2

@@ 3.2. Existence for approximated solutions
$\leq$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t$
$-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon,k}\cdot\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\rvert^{2}\,{\rm d}x\,{\rm d}t$
$\leq$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t$
$+\frac{\sigma^{2}}{2\varepsilon}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\right]^{2}|\nabla\rho_{\varepsilon,k}|^{2}$
$-\frac{\varepsilon}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\rvert^{2}\,{\rm d}x\,{\rm d}t,$ (20)
@@
$-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t\leq 0.$
The second one is that $\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}$ is Lipschitz continuous with respect to $(z_{n})_{+}$. In fact, we have the following
@@
$\left(\int_{\mathbb{R}^{d}}\rvert\rho_{\varepsilon,n}-\rho_{\varepsilon,k}\lvert^{2}\,{\rm d}x\right)(t)\leq$ $\;\frac{\chi\sigma^{2}}{2\varepsilon}\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert(z_{n})_{+}-(z_{k})_{+}\rvert^{2}\,{\rm d}x\,{\rm d}t$
$-\frac{\varepsilon}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\rvert^{2}\,{\rm d}x\,{\rm d}t.$
We conclude that $\{\nabla\rho_{\varepsilon,n}\}$ is a Cauchy sequence in $L^{2}(0,T,L^{2}(\mathbb{R}^{d})))$, and therefore, $\nabla\rho_{\varepsilon,n}\rightarrow\nabla\rho_{\varepsilon}$ a.e. $x\in[0,T]\times\mathbb{R}^{d}$. The map $\mathcal{T}$ is continuous, compact and constant when $\sigma=0$. Moreover, fixed points are uniformly bounded in $\sigma$.
Finally, we have to prove that fixed points of $\mathcal{T}$ are nonnegatives so we can remove the $+$ sign. Using Stampacchia’s truncation argument, we introduce $(\rho_{\varepsilon})_{-}$ in (11) as test function and we get
@@
If $\rho_{{\rm in},\varepsilon}\geq 0$, solutions to $\mathcal{T}(\sigma,\rho_{\varepsilon})=\rho_{\varepsilon}$ are nonnegatives.
Then, Leray-Schauder fixed-point yields a solution to
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t,$ (21)
for $\forall\;\varphi\in L^{2}(0,T,H^{1}(\mathbb{R}^{d})$.
4. The limit $\varepsilon\rightarrow 0$ and proof of Theorem 1
The main difficulty in removing the viscosity terms $\varepsilon\Delta\rho_{\varepsilon}$ and $-\varepsilon\rho_{\varepsilon}$ is to pass to the limit in the nonlinear term $\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}$. We need to show that
@@ 4. The limit $\varepsilon\rightarrow 0$ and proof of Theorem 1
We already know that
@@
The identification $G=\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)_{+}\nabla\rho$ will follow by almost everywhere convergence of $\rho_{\varepsilon}$ (given by Lemma (4)) and of $\nabla\rho_{\varepsilon}$. The next two subsections will be devoted to proving that
$\nabla\rho_{\varepsilon}\rightarrow\nabla\rho\quad\mbox{a.e. in }\;{((0,T)\times\mathbb{R}^{d})\backslash\{|\nabla\rho|\leq\chi\rho\}.}$
Then, in the last two subsections, we will prove the $L^{p}$ convergence of $\rho$ to initial data $\rho_{{\rm in}}$ and the uniqueness of the solutions.
4.1. Entropy method
The major tool of this argument is the relative entropy between two functions $u$ and $v$, defined as:
@@ 4.1. Entropy method
and $H[u|v]=0$ if and only if $u=v$. Let us outline how we use (22) to show that
Consider $\rho_{\varepsilon}$ and $\rho_{\delta}$ two viscosity solutions to (3.2) with viscosity constants $\varepsilon$ and $\delta$ respectively. A formal computation shows that
$H[\rho_{\varepsilon}|\rho_{\delta}](T)-H[\rho_{{\rm in},\varepsilon}|\rho_{{\rm in},\delta}]+\int_{0}^{T}\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]\,{\rm d}t=\mbox{error terms.}$. (23)
The relative entropy production $\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]$ is nonnegative as a direct consequence of the monotonicity of the operator $\left(1-\frac{\chi}{|\nabla\log\rho|}\right)_{+}\nabla\log\rho$. The error terms originate from the viscosities $\varepsilon\Delta\rho_{\varepsilon}$ and $\delta\Delta\rho_{\delta}$. As $\varepsilon,\delta\rightarrow 0$, $H[\rho_{{\rm in},\varepsilon}|\rho_{{\rm in},\delta}]\rightarrow 0$ by construction, and the error terms also vanish. Since both $H$ and $\mathcal{D}$ are nonnegative, as $\delta,\varepsilon\to 0$, we must have
@@
we obtain the almost everywhere convergence of $\nabla\rho_{\varepsilon}$ in the nondegenerate region $\{|\nabla\rho|>\chi\rho\}$. Consequently, we can identify the limit $\left(1-\frac{\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}$.
@@
One can make this argument rigorous and obtain (23) by testing (3.2) for $\rho_{\varepsilon}$ and for $\rho_{\delta}$ respectively with test functions $\varphi_{1}=\log\left(\frac{\rho_{\varepsilon}}{\rho_{\delta}}\right)$ and $\varphi_{2}=\frac{\rho_{\varepsilon}}{\rho_{\delta}}$. Since neither $\varphi_{1}$ nor $\varphi_{2}$ are admissible test functions, we modify them accordingly, at the expense of some error terms. Specifically, our test functions are of the form
@@
$\varphi_{1}=\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R},\;\;\mbox{and }\;\varphi_{2}=\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\phi_{R},$
with $\sigma>0$ a small constant and $\phi_{R}$ a cut-off function supported in $B_{R}$. More precisely, we chose $R=\sigma^{-\frac{1}{\alpha d}}$, with $\alpha>0$ to be fixed, and $\phi_{R}\in C^{\infty}_{c}(B_{2R})$ such that $\phi_{R}\equiv 1$ in $B_{R}$ and $|\nabla\phi_{R}|\leq\frac{C}{R}$ for some $C>0$. With this choice of $\varphi_{1}$ and $\varphi_{2}$ we obtain
$\int_{\mathbb{R}^{d}}$ $\left[\varphi_{1}\partial_{t}\rho_{\varepsilon}-\varphi_{2}\partial_{t}\rho_{\delta}\right]\,{\rm d}x=$
$=\frac{d}{dt}\int_{\mathbb{R}^{d}}\left[(\rho_{\varepsilon}+\sigma)\log{\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)}-\rho_{\varepsilon}+\rho_{\delta}\right]{{\phi_{R}}}\,{\rm d}x-\frac{d}{dt}\int_{\mathbb{R}^{d}}\rho_{\delta}\phi_{R}\,{\rm d}x.$ (24)
By Lemma 4, we know that
Hence, we can chose $\varepsilon,\delta>0$ small enough such that
$\int_{0}^{T}\int_{\mathbb{R}^{d}}|\rho_{\varepsilon}-\rho_{\delta}|^{2}\,{\rm d}x\,{\rm d}t<\sigma^{2}.$ (25)
We do not write $\varepsilon(\sigma),\delta(\sigma)$ and $R(\sigma)$ on $\sigma$ for readability purposes, but notice that
$\lim_{\sigma\rightarrow 0}\varepsilon,\delta=0,\quad\textrm{and}\quad\lim_{\sigma\rightarrow 0}R(\sigma)=\infty.$
Next, we subtract equation (3.2) for $\rho_{\delta}$, tested with $\varphi_{2}$, to the equation (3.2) for $\rho_{\varepsilon}$, tested with $\varphi_{1}$. This yields:
@@
$\int_{0}^{T}$ $\int_{\mathbb{R}^{d}}\left[\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\partial_{t}\rho_{\varepsilon}-\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\;\partial_{t}\rho_{\delta}\right]\phi_{R}\,{\rm d}x\,{\rm d}t$ (26)
$=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\left[\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}+\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\nabla\phi_{R}\right]\,{\rm d}x\,{\rm d}t$
$+\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\left[\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}+\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\nabla\phi_{R}\right]\,{\rm d}x\,{\rm d}t+E_{\varepsilon,\delta},$
where $E_{\varepsilon,\delta}$ are error terms, which must vanish when $\sigma\rightarrow 0$. Notice that (26) is an approximation of $\int_{0}^{T}\frac{d}{\,{\rm d}t}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}]$. We will show
$\lim\limits_{\varepsilon,\delta\rightarrow 0}\int_{0}^{T}\frac{d}{\,{\rm d}t}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}]\leq 0,$
to find that, in fact, $\lim\limits_{\varepsilon,\delta\rightarrow 0}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}](T)=0$. This limit will provide the necessary information about the convergence of $\nabla\rho_{\varepsilon}$.
To do this, let us analyze the error terms first:
$E_{\varepsilon,\delta}=$ $-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t+\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\delta}\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t$ (27a)
@@
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\phi_{R}\left[\varepsilon\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)-\delta\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right]\,{\rm d}x\,{\rm d}t$ (27b)
@@
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\phi_{R}\left[\varepsilon\nabla\rho_{\varepsilon}\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)-\delta\nabla\rho_{\delta}\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right]\,{\rm d}x\,{\rm d}t.$ (27c)
@@
For the first term above, assuming without loss of generality $\varepsilon<\delta$ and recalling that we can always choose $\delta$ such that $\delta<\sigma<<1$ we compute

### v3

@@ 3.2. The nonlinear problem
$\leq$ $-\tilde{\sigma}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t$
$-\tilde{\sigma}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\hskip-5.69054pt-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon,k}\cdot\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\lvert\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\rvert^{2}\,{\rm d}x\,{\rm d}t$
$\leq$ $-\tilde{\sigma}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t$
$+\frac{\tilde{\sigma}^{2}}{2\varepsilon}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\hskip-5.69054pt-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\right]^{2}|\nabla\rho_{\varepsilon,k}|^{2}\,{\rm d}x\,{\rm d}t$
$-\frac{\varepsilon}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\lvert\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\rvert^{2}\,{\rm d}x\,{\rm d}t,$ (23)
@@
$-\tilde{\sigma}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t\leq 0.$
[…]
@@
$\left(\int_{\mathbb{R}^{d}}\rvert\rho_{\varepsilon,n}-\rho_{\varepsilon,k}\lvert^{2}\,{\rm d}x\right)(t)\leq$ $\;\frac{\chi\tilde{\sigma}^{2}}{2\varepsilon}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\lvert(z_{n})_{+}-(z_{k})_{+}\rvert^{2}\,{\rm d}x\,{\rm d}t$
$-\frac{\varepsilon}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\lvert\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\rvert^{2}\,{\rm d}x\,{\rm d}t.$
We conclude that $\{\nabla\rho_{\varepsilon,n}\}$ is a Cauchy sequence in $L^{2}(0,T,L^{2}(\mathbb{R}^{d})))$, and therefore, $\nabla\rho_{\varepsilon,n}\rightarrow\nabla\rho_{\varepsilon}$ a.e. $x\in[0,T]\times\mathbb{R}^{d}$.
Summarizing, we have shown that the map $\mathcal{T}$ is continuous, compact and constant when $\tilde{\sigma}=0$. Moreover, fixed points are uniformly bounded in $\tilde{\sigma}$. Then, Leray-Schauder fixed-point theorem ensure the existence of $\rho_{\varepsilon}$ such that $\mathcal{T}(1,\rho_{\varepsilon})=\rho_{\varepsilon}$, which is the solution to the nonlinear problem
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi{\rho_{\varepsilon}}_{+}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t.$
To conclude the proof of existence of solutions to (19), it remains to show that $\rho_{\varepsilon}$ is nonnegative. For that, use $(\rho_{\varepsilon})_{-}$ in the above weak formulation and obtain
@@
Since $\rho_{{\rm in},\varepsilon}\geq 0$ the last inequality implies ${\rho_{\varepsilon}}_{-}\equiv 0$. This concludes the proof.
4. The vanishing viscosity limit and proof of the main theorem
The aim of this section is to remove the viscosity terms $\varepsilon\Delta\rho_{\varepsilon}$ and $\varepsilon\rho_{\varepsilon}$ from (19). As explained in the introduction, the main difficulty is to pass to the limit in the nonlinear term $\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}$ and show that
@@ 4. The vanishing viscosity limit and proof of the main theorem
From Proposition 1 we know that
@@
Moreover, Lemma (4) implies $\rho_{\varepsilon}\to\rho$ a.e.. The crux is to show that $G=\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)_{+}\nabla\rho$.
Using the Boltzmann relative entropy we will show that $\nabla\rho_{\varepsilon}$ converges a.e. to $\nabla\rho$; in particular we will prove that
[…]
Let’s briefly sketch how we get the last a.e. convergence; the relative entropy between two functions $u$ and $v$ is defined as:
@@
and $H[u|v]=0$ if and only if $u=v$. Consider $\rho_{\varepsilon}$ and $\rho_{\delta}$ two viscosity solutions to (19) with viscosity constants $\varepsilon$ and $\delta$ respectively. A formal computation shows that
$H[\rho_{\varepsilon}|\rho_{\delta}](T)-H[\rho_{{\rm in},\varepsilon}|\rho_{{\rm in},\delta}]+\int_{0}^{T}\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]\,{\rm d}t=\mbox{error terms.}$ (25)
It turns out that the relative entropy production $\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]$ is nonnegative, a direct consequence of the monotonicity property (10). The error terms originate from the viscosities $\varepsilon\Delta\rho_{\varepsilon}$ and $\delta\Delta\rho_{\delta}$. As $\varepsilon,\delta\rightarrow 0$ these error terms vanish, as well as $H[\rho_{{\rm in},\varepsilon}|\rho_{{\rm in},\delta}]\rightarrow 0$. Since both $H$ and $\mathcal{D}$ are nonnegative, as $\delta,\varepsilon\to 0$, we must have
@@
we will obtain the almost everywhere convergence of $\nabla\rho_{\varepsilon}$ to $\nabla\rho$ in the nondegenerate region $\{|\nabla\rho|>\chi\rho\}$ and consequently will be able to identify the limit function $G$ as $G=\left(1-\frac{\rho}{|\nabla\rho|}\right)_{+}\nabla\rho$. The rigorous proof of this argument is the content of sections 4.1 and 4.2.
Then, in the last two subsections, we will prove the $L^{p}$ convergence of $\rho$ to initial data $\rho_{{\rm in}}$ and the uniqueness.
4.1. Contractivity of the relative entropy
@@
In what follows, we will make the argument above rigorous. We start by testing (19) for $\rho_{\varepsilon}$ and for $\rho_{\delta}$ respectively with test functions $\varphi_{1}=\log\left(\frac{\rho_{\varepsilon}}{\rho_{\delta}}\right)$ and $\varphi_{2}=\frac{\rho_{\varepsilon}}{\rho_{\delta}}-1$ (this will formally give $\frac{d}{dt}H[\rho_{\varepsilon}|\rho_{\delta}]$). Since neither $\varphi_{1}$ nor $\varphi_{2}$ are admissible test functions, we modify them accordingly, at the expense of some error terms. Specifically, our test functions are of the form
@@
$\varphi_{1}=\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right),\;\;\mbox{and }\;\varphi_{2}=\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}-1,$
with $\sigma>0$ a small constant. We point out that $\varphi_{1}$, $\varphi_{2}$ are indeed admissible test functions as they belong to $L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$. For example,
$|\varphi_{1}|\leq\log(1+\rho_{\varepsilon}/\sigma)+\log(1+\rho_{\delta}/\sigma)\leq(\rho_{\varepsilon}+\rho_{\delta})/\sigma\in L^{\infty}(0,T;L^{1}\cap L^{\infty}(\mathbb{R}^{d})).$
The integrability of $\varphi_{2}$, $\nabla\varphi_{1}$, $\nabla\varphi_{2}$ can be easily checked.
With this choice of $\varphi_{1}$ and $\varphi_{2}$ we obtain
@@
$\int_{\mathbb{R}^{d}}$ $\left[\varphi_{1}\partial_{t}\rho_{\varepsilon}-\varphi_{2}\partial_{t}\rho_{\delta}\right]\,{\rm d}x=\frac{d}{dt}\int_{\mathbb{R}^{d}}\left[(\rho_{\varepsilon}+\sigma)\log{\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)}-\rho_{\varepsilon}+\rho_{\delta}\right]\,{\rm d}x.$ (26)
This yields:
@@
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\varphi_{1}\partial_{t}\rho_{\varepsilon}-\varphi_{2}\partial_{t}\rho_{\delta}\right]\,{\rm d}x\,{\rm d}t=$ $\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\partial_{t}\rho_{\varepsilon}+\left(1-\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\;\partial_{t}\rho_{\delta}\right]\,{\rm d}x\,{\rm d}t$ (27)
@@
$=$ $-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\cdot\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t$
$+\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\cdot\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t+E_{\varepsilon,\delta,\sigma},$
where $E_{\varepsilon,\delta,\sigma}$ are error terms, which must vanish when $\varepsilon,\delta,\sigma\rightarrow 0$. Notice that (26) is an approximation of $\int_{0}^{T}\frac{d}{\,{\rm d}t}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}]\;dt$. We will show
$\lim\limits_{\varepsilon,\delta\rightarrow 0}\int_{0}^{T}\left(\frac{d}{\,{\rm d}t}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}]{+\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]}\right)\,{\rm d}t\leq 0,$
to find that, in fact, $\lim\limits_{\varepsilon,\delta\rightarrow 0}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}](T)=0$ and $\lim\limits_{\varepsilon,\delta\rightarrow 0}\int_{0}^{T}\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]\,{\rm d}t=0$. These limits will provide the necessary information about the convergence of $\nabla\rho_{\varepsilon}$.
Let us first analyze the error terms:
$E_{\varepsilon,\delta,\sigma}=$ $-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\,{\rm d}x\,{\rm d}t+\delta\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\delta}{\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}-1\right)}\,{\rm d}x\,{\rm d}t$ (28a)
$-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\varepsilon\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)-\delta\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right]\,{\rm d}x\,{\rm d}t.$ (28b)
We notice first that
$\left|\sigma\log\left(\frac{a+\sigma}{b+\sigma}\right)\right|=\left|\sigma\log\left(1+\frac{a}{\sigma}\right)-\sigma\log\left(1+\frac{b}{\sigma}\right)\right|\leq a+b\quad\forall a,b\geq 0.$ (29)
@@
For the first term in $E_{\varepsilon,\delta}$, assuming without loss of generality $\varepsilon<\delta<\frac{1}{2}$ we compute (remember that $H_{rel}[\rho_{\varepsilon}|\rho_{\delta}]\geq 0$)